OpenStudy (anonymous):

find the least degree of a polynomial equation with roots -1, 3, and +-3i I have no idea how to solve this. Can anyone help explain it to me?

OpenStudy (ranga):

There are four roots and therefore the least degree will be 4 (That is x^4 + .... ) If a is a root it means (x-a) is a factor. So to get the polynomial, just multiply: (x+1)(x-3)(x-3i)(x+3i) Multiply the last two factors first to get rid of the i. Also, make use of the identity: (a+b)(a-b) = a^2 - b^2 when multiplying the last two factors.

OpenStudy (anonymous):

so (x+1)(x-3)(x^2-9)?

OpenStudy (ranga):

I^2 = -1 and so try again.

OpenStudy (anonymous):

would the last term be x^2+9 then?

OpenStudy (anonymous):

(x-3i)(x+3i) foil it x times x is x^2 -3i times 3i is -3i^2? or would you break it out into -3 times 3 and -i times i either way -i times i is 1, because i times i is -1 right

OpenStudy (ranga):

(-3i) * (3i) = -9(i^2) = -9(-1) = +9 (x-3i)(x+3i) = x^2 + 9 Multiply that by (x+1) and then by (x-3). Arrange the polynomial from the highest degree to the lowest degree.